# Toward the Categorical Data Map

### v3

@@ VII Conclusion
We presented a novel projection-based visualization method to address the need for similarity-based analysis techniques for categorical data by leveraging distance relations based on set intersections to create enhanced, interactive glyph-based scatterplot-like visualizations called the Categorical Data Map. Because of the inherent sparsity of categorical data, groups of similar data items are grouped into clusters. We visualized attributes and categories by calculating a Voronoi partitioning and coloring the cells according to the category of the associated attribute. Our method allows for exploring the categorical data space through segmentation, enabling the orientation along an automatic or user-selected attribute. For automatic selection, we rank-order attributes along a visual property we defined as fracturedness, for which we provided two measures. Through a case study, we showed that our Categorical Data Map can support the identification of similar subsets and clusters, as well as the detection of attributes with a strong influence on the topology of the embedding. Additionally, in an expert study, we were able to confirm that our approach facilitates the analysis of categorical data, especially for large datasets, by grouping similar subsets while, through layout enrichment, visualizing the distribution of categories of an attribute. We published a demonstrator and our results online so that users can interactively experiment with our approach and build upon our results. We conclude that the Categorical Data Map provides an effective way to analyze large categorical datasets, especially in exploratory scenarios.

### v4

@@ 6 Conclusion
We presented a novel projection-based visualization method to address the need for similarity-based analysis techniques for categorical data. We leverage distance relations based on set intersections to create enhanced and interactive glyph-based scatterplot-like visualizations called the Categorical Data Map. We visualized attributes and categories by calculating a Voronoi partitioning and coloring the cells according to the category of the associated attribute. Our method allows for exploring the categorical data space through segmentation, enabling the orientation along an automatic or user-selected attribute. For automatic selection, we rank-order attributes along a visual property we defined as fracturedness measures. We quantitively evaluated different distance measures for the projection of categorical data with MDS, suggesting that the Overlap coefficient and Jaccard distance yield results outperforming MCA. Through a case study, we showed that our Categorical Data Map can support the identification of similar subsets and clusters, as well as the detection of attributes with a strong influence on the topology of the embedding. In an expert study, we were able to confirm that our approach facilitates the analysis of categorical data, especially for large datasets, by grouping similar subsets while, through layout enrichment, visualizing the distribution of categories of an attribute. We published a demonstrator and our results online so that users can interactively experiment with our approach and build upon our results. We conclude that the Categorical Data Map effectively analyzes large categorical datasets, especially in exploratory scenarios.
Acknowledgements.
We thank the anonymous reviewers for their valuable feedback. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 251654672 – TRR 161 (Projects A01 and A03) and the Federal Ministry for Economic Affairs and Climate Action (BMWK, grant No. 03EI1048D).